OpenStudy (anonymous):

Currently, Sam's coach is three times older than Sam. When Sam started soccer 5 years ago, his coach was five times his age. How old is Sam now?

OpenStudy (anonymous):

sam - s coach - c c-5=5(s-5) c=3s solving these two we get s=10 c=30